LUCIFER Training & Preparations for Partner Science:
UT 2010 May 01

Observer: DThompson (LBTO)
Telescope Operator: WWack
Instrument Support: TShih
Telescope Support: JLittle, JUrban
SW support: DCox(on call)

Summary

  • Telescope started out warm (+10C), thermalization took a while and made seeing worse
  • Repeated (4x) problems with OSS dropping out unannounced
  • Collected some questionable partner data:
    • OSU: UM425 spectra + calibrations
    • INAF: sn2010aj JHK imaging
  • Remeasured IA after telescope thermalized, IA=-400. Updated pointing model + defaults
  • Collected some equally questionable LBTO data:
    • Two B3.2 sources, Ks imaging for Marco
    • Test spectra on HR10, DThompson
    • Could not do XRB/Wagner, moon too close to the targets to collimate with selected G*
  • Some data taken with archive section in scripts

  • Notes for LUCIFER Team:
    • Script processing xterm shows offsets in RA/DEC even when they are done in DETX/Y
    • OBJECT field on RMGUI remains "NOTSPECIFIED" in all conditions. How is this set?
    • Skycat never auto-displayed images, despite a restart with ~/bin/start_skycat.sh
    • Can FLUSH_DETECTOR be made configurable? That is: "FLUSH_DETECTOR = 150" for 150 reads.

Details

02:10 It has mostly cleared up. M1 is close to 10C warmer than outside so we are ventilating and the vent doors are open but the main doors are still closed.

04:05 Opening enclosure. Sky looks clear, RH=47%. Tout=-7.9, Tmirror=-1.8 (M1 is 6.1C warm)

04:20 Presets failing for M3. Called Paul, talked Wayne through setting M3 left in simulation mode. Subsequent preset succeeded.

05:00 OSS died? Collimation updates do not look like they are being applied. Restarting OSS.

05:15 Seeing is ~2" but running OSUdietrich.UM425_acq anyway for practice.
  • 0006 through-slit acquisition
  • 0007 image, offset 5,5
  • 0008 image, back to starting pos
  • 0009 source through slit image...in slit but M1 hit collimation limit!

05:38 re-execute acq script. OSS dropped out on collimating M2. Kill and restart OSS at 05:43. Skycat was not displaying images so I restarted skycat from the ~/bin directory.

05:54 Preset to BS9149 after correcting pointing there, allow to collimate.

06:05 re-execute UM425 acq script.
  • 0013 through slit acq image
  • 0014 on sky image
  • 0015 source on slit. Shift image worked. Skycat still not auto-displaying data.

06:16 execute UM425 Science script. OSS hung again. Exit script. Calling Dan Cox. While he looks into things...

06:23 re-execute UM425 acq script.
  • 0016 through slit image
  • 0017 through slit image after applying 4.73, 2.08 offsets from last shift image...worked!

06:26 execute UM425 sci script. Guider showing 1.5 arcsec FWHM. Files: 0018-0023.

06:51 execute HIP54849 acq script...failed (OSS dropped out again).
07:00 re-execute...preset failed, no guide star. Correct pointing...
  • 0025 through slit
  • 0026 source image

07:15 execute HIP54849 science script. Files 0027-0030, plus arcs and flats with calibration unit.

07:35 execute Turatto sn2010al by accident (EL=2.0!).

07:39 Wait for collimation was off. Exit script, correct pointing (star was at edge) and resubmit.

07:45 execute Turatto sn2010aj Ks imaging (EL~36 at start), files 0053-0062.
08:04 execute Turatto sn2010aj J imaging (EL~33 at start), files 0063-0067.
08:14 execute Turatto sn2010aj H imaging (EL~31 at start), first preset failed, homing AGw stages (scripts have same preset info, no other reaosn to fail). Re-execute at 08:18UT, files 0068-0072. At least the first couple look elongated, as do the guider images.
These data are likely not dithered enough! The galaxy overlaps itself in all dithers, so there will be no clean sky subtraction.

08:30 Pointing correction near HR10. Collimation successful 08:39:28
  • 0074 sky image
  • 0075 on-source image (at preset position) (not dithered...REL offset of 0 0)
  • 0075 on-source image

  • 0079-0080 300s H spectra. Nothing obvioius. Check exposure time vs. sky lines (brightest is 5000 counts in 300s)
  • 0081-0085 Ks imaging 6x10s to adjust position angle?

09:23 Pointing to B3.2 1532+0054
09:28 execute Ks_600s.script on this field, files 0086-0095

09:45 Correct pointing nearby, 0096

09:47 Pointing to B3.2 1534+0001, execute Ks_600s.script, files 0097-0106

10:06 Correct pointing near Mark's field

10:10 Slit07_acq executed. Star is faint and there is a lot of background from the moon on the WFS...not able to collimate because of the scattered light from the moon. Field is identifiable in decent seeing.

10:22 Check IA term again... IA=-400. Modified PCSInstrument.conf to use LUCIFERSX.20100501.ptmod by default. This is the 20100103 file with IA fixed at -400.

10:53 Practice alignment on slit more on HR10 field.
  • 0115 sky
  • 0116 image
  • 0117 through-slit

12:00 Closing dome. Guide star lost.

DoDarks:

-- DavidThompson - 01 May 2010
Topic revision: r8 - 05 May 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback